The pieces of a 500 piece puzzle are stored in three containers. 220 pieces are in the first container and 180 pieces are in the second container. what percentage of the pieces is in the third container?
Answer by addingup(3677)

500-220-180 = 100
(100/500)*100 = % of pieces in the third container